# Associate Everything Detected: Facilitating Tracking-by-Detection to the Unknown

### v1

@@ III-C Association-Centric Learning Mechanism
To better understand our approach, we visualize the association-centric learning in Fig. 4. The top-left panel of Fig. 4 provides an example of a sampled clip containing three frames during training. Assuming that there are four IDs in this clip, ID 1 and ID 4 are absent in frame 1 and frame 3 respectively. During training, their corresponding object queries are encoded from real detections.
For each $W$ output by sim-decoder, we apply the embedding loss proposed in [22] to enhance the weight responses among keys and queries that share the same ID and diminish the responses from different IDs. We formulate the loss as follows:
@@
where $r$ is the index of rows in $W$, and each row of $W$ represents the responses of all keys to the same query in multi-head weight attention. $w^{+}_{r},w^{-}_{r}\in W_{r}$ and $W_{r}$ denotes the $r$-th row of $W$. $w^{+}_{r}$ is the response between key and query from the same ID, and $w^{-}_{r}$ is the response between different IDs. We further adopt focal loss [39] as the auxiliary loss based on the corresponding $S$:
@@
We argue that distinguishing different IDs within a single frame is fundamental for a tracker before cross-frame data association. Additionally, we all know that objects within a single frame are unique. Thus, we leverage $W_{\mathrm{spatial}}$ and $S_{\mathrm{spatial}}$ to enhance the spatial discriminative ability of AED. Specifically, we use $\mathcal{L}_{\mathrm{embed}}$ and $\mathcal{L}_{\mathrm{aux}}$ to accomplish spatial contrastive learning:
@@
Based on spatial contrastive learning, temporal contrastive learning aims to connect object queries with track queries that share the same ID. Again, we perform matching to achieve real track queries during training. Thus, by using $W_{\mathrm{temporal}}$ and $S_{\mathrm{temporal}}$, temporal contrastive learning can be formulated as:
@@
Since the ultimate object of MOT is to keep the ID as long as possible, we further use cross-clip contrastive learning to further intensify the long-term discriminability of AED using $W_{\mathrm{cross-clip}}$ and $S_{\mathrm{cross-clip}}$. Similarly, for the entire clip, we define the cross-clip contrastive learning as:
@@
To sum up, association-centric learning aligns the training process with the association process making AED more robust to different tracking scenarios.
@@ III-D Training Details
$+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}$
@@
We use real detected boxes to generate object queries during training including the boxes that are not matched with any GT, e.g. false positives. The matched boxes (i.e. IoU with GT $>0.5$) will contribute to all of the loss functions. The unmatched boxes will contribute to $\mathcal{L}_{\mathrm{spatial}}$ and will act as negative samples in $\mathcal{L}_{\mathrm{temporal}}$. If any object query is matched with a GT initially but its IoU with GT is smaller than a threshold $\tau_{\mathrm{iou}}$ after box refinement, we simply filter it out to avoid ambiguous learning.
For OV-MOT tasks, we split categories into $\mathcal{C}^{\mathrm{base}}$ and $\mathcal{C}^{\mathrm{novel}}$ following [13]. We only use $\mathcal{C}^{\mathrm{base}}$ to train AED, and we use $\mathcal{C}^{\mathrm{base}}$ and $\mathcal{C}^{\mathrm{novel}}$ to evaluate our models. For CV-MOT tasks, we utilize the entirety of the annotated data for training.
@@ IV Experiments
TABLE I: Comparison with state-of-the-art OV-MOT trackers on TAO dataset. We categorize the experimental results based on the differences in detectors. ”Intergrated” indicates this method uses its model to localize and extract features of objects simultaneously. The best results are shown in bold and the second best results are shown with underlines.
@@ IV-A Datasets And Evaluation Metrics
Datasets. We use TAO [34], DanceTrack [9], and SportsMOT [25] datasets to evaluate both OV-MOT and CV-MOT performance of AED. TAO is a large-scale dataset that contains 482 categories for MOT tasks. It is annotated under the federated protocol [44], and there are up to 10 labeled objects in every video of TAO. Additionally, TAO is annotated at a quite low frame rate, i.e. 1 FPS. SportsMOT aims to track players on the court, including basketball, volleyball, and football. Except for similar appearances, SportsMOT also has fast and variable-speed motion, which poses a higher challenge to the performance of the tracker. DanceTrack also has one category: person. It is a challenging dataset with diverse motion and severe occlusion, and people in DanceTrack also have very similar appearances.
@@ IV-B Implementation Details
We assign each detection box to a GT box based on an IoU threshold ($>0.5$) during the clip sampling process. Since the TAO dataset is annotated under the federated protocol [44], i.e., not all instances of the same category are annotated in the images, the treatment of these unmatched boxes follows Section III-D.
@@ IV-C Comparison With OV-MOT Trackers
In Tab. I, we give the evaluation result on both validation and test sets of TAO following [13]. AED is trained on $\mathcal{C}^{\mathrm{base}}$ categories on TAO and the object queries are generated from Co-DETR [35]. During inference, we use two off-the-shelf detectors, RegionCLIP (RN50 backbone version) [11] and Co-DETR [35], to provide bounding boxes for AED. Due to the categories of the TAO dataset are consistent with those of LVIS, RegionCLIP and Co-DETR are both trained on LVIS [44].
When using RegionCLIP as the detector, AED achieves outstanding performance under the TETA metric compared with several high-performance methods [22, 42, 13, 15, 2, 16, 43]. The previous state-of-the-art method OVTrack[13] uses a dual contrastive training strategy and achieves excellent performance. AED outperforms OVTrack by 2.5% (36.3%-37.2%) on TETA of base categories and 5.6% (32.0%-33.8%) on TETA of the unseen novel categories in the validation set of TAO. When it comes to the test set, AED also has an increment of 6.9% (34.8%-37.2%) on TETA of base categories and 8.2% (25.7%-27.8%) on TETA of the novel categories in the validation set of TAO. Since AED is responsible for associating every detected bounding box, the AssocA metric directly reflects the performance of AED. AED outperforms OVTrack by 4.4% (36.3%-37.9%) and 12.0% (33.2%-37.2%) on AssocA of base and novel categories of validation set respectively. For the test set, the increment for AssocA on base and novel categories are 11.9% (36.1%-40.4%) and 11.1% (26.2%-29.1%) respectively.
@@ IV-D1 SportsMOT
We further validate the CV-MOT performance of AED on the SportsMOT dataset. As shown in Tab II, for AED, we first use the weights trained on the $\mathcal{C}^{\mathrm{base}}$ of TAO following Section IV-C while the detector (YOLOX [36]) is trained on the training set of SportsMOT. Under such a setting, AED achieves a competitive HOTA score, i.e. 72.8, outperforming a large number of trackers [50, 22, 51, 5, 52, 16, 17, 53]. For a more equitable comparison, we also train AED, which is pretrained on COCO, on the training set of SportsMOT. AED achieves state-of-the-art performance on most metrics. Compared with Deep-EIoU [23], AED achieves the increment of 3.9% (74.1%-77.0%) on HOTA and 6.7% (75.0%-80.0%) on IDF1 by using only training set. Note that unlike other tracking-by-detection methods leveraging some extra prior knowledge, AED only uses the similarities of features to achieve this performance.
@@ IV-E1 Analysis on Association-Centric Learning Mechanism
TABLE VI: Multi-head weight attention ablation on the validation set of TAO dataset.
@@
IV-E2 Weight Attention Analysis
In Tab. VI, we conduct several experiments to validate the impact of different $h$ in the multi-head weight attention. We select several different values of $h$, i.e., 1, 2, 4, 8, and 16. It can be observed from Tab. VI that when $h$ equals 2, the model achieves a comparatively high performance.
TABLE VII: $\tau_{iou}$ ablation on the validation set of TAO dataset.
@@ IV-E2 Weight Attention Analysis
IV-E3 Analysis on IoU Threshold to Filter Out Ambiguous Quries
As introduced in Section III-D, after the refinement in training, we calculate IoUs between boxes of object queries and their GTs to filter out the ambiguous ones, i.e. IoU $<\tau_{\mathrm{iou}}$. In order to investigate the impact of $\tau_{\mathrm{iou}}$ on AED, we conduct a comparative experiment in Tab. VII. We find that a higher $\tau_{\mathrm{iou}}$ has a slight negative impact on the association performance (AssocA) of AED. Ultimately, we chose $\tau_{\mathrm{iou}}=0.5$ during training.
TABLE VIII: The impact of different clip lengths on the performance of AED.
@@ IV-E3 Analysis on IoU Threshold to Filter Out Ambiguous Quries
IV-E4 Analysis on the Length of Clips
During training, we take sampled clips of $n$ frames as the input. We conduct an experiment on the TAO dataset to find the most suitable $n$ in Tab. VIII. We can find that with the increment of $n$, the performance of AED also improves. However, due to the limitation of GPU memory, $n=5$ represents the current upper limit of our experimental conditions.
IV-E5 Analysis on Sampling Strategy
As the input of AED is a video clip in the training stage, there are different ways to merge a number of video frames into a clip. We conduct ablation experiments on two sampling methods on the TAO dataset as shown in Tab. IX. The first setting is that we fix the sampling interval to the number of three, i.e., skip two frames when sampling. Second, we allowed the sampling interval a to change randomly within the range of [0, 3], which introduces a wider variety of displacement and perspective changes during training. We find that the random sampling strategy improves the AssocA by 0.6 points.
TABLE IX: Different sampling strategy ablation on the validation set of TAO dataset.
Sampling Methods TETA LocA AssocA ClsA
fixed 54.9 71.7 51.5 41.6
random 55.2 71.8 52.1 41.7
@@ IV-E5 Analysis on Sampling Strategy
IV-E6 Analysis on Performance Ceiling
IV-F Visualization
IV-F1 Association-Centric Learning
IV-F2 Tracking Results
In Fig. 6, we visualize some successful and failed tracking results of AED on different datasets including TAO, SportsMOT, and DanceTrack. Specifically, we use Co-DETR in TAO and YOLOX in SportsMOT and DanceTrack. AED is able to handle a wide range of scenarios including low frame rates (TAO), large movements(SportsMOT), severe occlusions (DanceTrack), and similar appearances (SportsMOT and DanceTrack). However, AED makes some mistakes in cases involving tiny objects (ID9, 10, 12, etc of the first row of the TAO dataset) and highly dense scenes (ID 64, 65, etc of the second row of the TAO dataset).

### v2

@@ III-C Association-Centric Learning Mechanism
To better understand our approach, we visualize the association-centric learning in Fig. 4. Fig. 4(a) provides an example of a sampled clip containing three frames during training. Assuming that there are four IDs in this clip, ID 1 and ID 2 are absent in frame 3 and frame 2 respectively. ID 4 only appears in frame 3. During training, their corresponding object queries are encoded from real detections.
For each $W$ (i.e., $W_{\mathrm{spatial/temporal/cross-clip}}$) output by sim-decoder, we apply the embedding loss proposed in [24] to enhance the weight responses among keys and queries that share the same ID and diminish the responses from different IDs. We formulate the loss as follows:
@@
where $r$ is the index of rows in $W$, and each row of $W$ represents the responses of all keys to the same query in multi-head weight attention. $w^{+}_{r},w^{-}_{r}\in W_{r}$ and $W_{r}$ denotes the $r$-th row of $W$. $w^{+}_{r}$ is the response between key and query from the same ID, and $w^{-}_{r}$ is the response between different IDs. We further adopt focal loss [41] as the auxiliary loss based on the corresponding $S$ (i.e., $S_{\mathrm{spatial/temporal/cross-clip}}$):
@@
We argue that distinguishing different IDs within a single frame is fundamental for a tracker before cross-frame data association. Additionally, we all know that objects within a single frame are unique. As shown in Fig. 4(b), a target should only have a high response with itself within a single frame. Thus, we leverage $W_{\mathrm{spatial}}$ and $S_{\mathrm{spatial}}$ to enhance the spatial discriminative ability of AED. Specifically, we use $\mathcal{L}_{\mathrm{embed}}$ and $\mathcal{L}_{\mathrm{aux}}$ to accomplish spatial contrastive learning:
@@
Temporal contrastive learning aims to connect object queries with track queries that share the same ID. During training, we perform matching to generate real track queries, rather than relying on information from two joint frames. This can be seen in Fig. 4(c), where ID2 from frame 1 is also included. Thus, by using $W_{\mathrm{temporal}}$ and $S_{\mathrm{temporal}}$, temporal contrastive learning can be formulated as:
@@
Since the ultimate object of MOT is to keep the ID as long as possible, we further use cross-clip contrastive learning to intensify the long-term discriminability of AED. As shown in Fig. 4(d), we use all the objects in this clip to learn the feature consistency of targets across multiple frames. Similarly, for the entire clip, we define the cross-clip contrastive learning as:
@@
Instead of introducing an additional memory module [31, 32], our method integrates long-term modeling directly into the training process and achieves significant computational savings.
To sum up, association-centric learning aligns the training process with the association process, making AED more robust to different tracking scenarios.
@@ III-D Training Details
$+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}},$
@@
We use real detected boxes to generate object queries during training including the boxes that are not matched with any GT, e.g., false positives. The matched boxes (i.e., IoU with GT >0.5>0.5> 0.5) will contribute to all of the loss functions. The unmatched boxes will contribute to $\mathcal{L}_{\mathrm{spatial}}$ and will act as negative samples in $\mathcal{L}_{\mathrm{temporal}}$. If any object query is matched with a GT initially but its IoU with GT is smaller than a threshold $\tau_{\mathrm{iou}}$ after box refinement, we simply filter it out to avoid ambiguous learning.
For OV-MOT tasks, we split categories into $\mathcal{C}^{\mathrm{base}}$ and $\mathcal{C}^{\mathrm{novel}}$ following [14]. Specifically, $\mathcal{C}^{\mathrm{novel}}$ are the rare classes defined in [44] and $\mathcal{C}^{\mathrm{base}}$ are the rest of the classes in TAO. We only use $\mathcal{C}^{\mathrm{base}}$ to train AED, and we use $\mathcal{C}^{\mathrm{base}}$ and $\mathcal{C}^{\mathrm{novel}}$ to evaluate our models. Introducing new categories during testing that are not seen in the training phase can evaluate the class generalization ability of the model. For CV-MOT tasks, we utilize the entirety of the annotated data for training.
@@ IV Experiments
TABLE I: Comparison with state-of-the-art OV-MOT trackers on TAO dataset. We categorize the experimental results based on the differences in detectors. ”Integrated” indicates this method uses its model to localize and extract features of objects simultaneously. The best results are shown in bold and the second best results are shown with underlines.
@@ IV-A Datasets And Evaluation Metrics
Datasets. We use TAO [36], DanceTrack [9], and SportsMOT [27] datasets to evaluate both OV-MOT and CV-MOT performance of AED. TAO is a large-scale dataset that contains 482 categories for MOT tasks. It is annotated under the federated protocol [44], and there are up to 10 labeled objects in every video of TAO. Additionally, TAO is annotated at a quite low frame rate, i.e., 1 FPS. SportsMOT aims to track players on the court, including basketball, volleyball, and football. Except for similar appearances, SportsMOT also has fast and variable-speed motion, which poses a higher challenge to the performance of the tracker. DanceTrack also has one category: person. It is a challenging dataset with diverse motion and severe occlusion, and people in DanceTrack also have very similar appearances.
@@ IV-B Implementation Details
We assign each detection box to a GT box based on an IoU threshold (>0.5>0.5> 0.5) during the clip sampling process. Since the TAO dataset is annotated under the federated protocol [44], i.e., not all instances of the same category are annotated in the images, the treatment of these unmatched boxes follows Section III-D.
@@ IV-C Comparison With OV-MOT Trackers
In Tab. I, we give the evaluation result on both validation and test sets of TAO following [14]. AED is trained on $\mathcal{C}^{\mathrm{base}}$ categories on TAO and the object queries are generated from Co-DETR [37]. During inference, we use two off-the-shelf detectors, RegionCLIP (RN50 backbone version) [12] and Co-DETR [37], to provide bounding boxes for AED. To ensure fair comparison, we use the same detector to evaluate different algorithms. Due to the categories of the TAO dataset are consistent with those of LVIS [44], RegionCLIP is trained on LVIS with $\mathcal{C}^{\mathrm{base}}$, and Co-DETR is trained on the full set of LVIS. We simply use a class-agnostic NMS (Non-Maximum Suppression) to remove duplicate detections.
When using RegionCLIP as the detector, AED achieves outstanding performance under the TETA metric compared with several high-performance methods [24, 45, 14, 16, 2, 18, 46]. The previous state-of-the-art method, OVTrack [14], uses a dual contrastive training strategy and achieves excellent performance. AED outperforms OVTrack by 2.8% (36.3%-37.3%) on TETA of base categories and 7.5% (32.0%-34.4%) on TETA of the unseen novel categories in the validation set of TAO. When it comes to the test set, AED also has an increment of 6.9% (34.8%-37.2%) on TETA of base categories and 8.2% (25.7%-27.8%) on TETA of the novel categories in the validation set of TAO. Since AED is responsible for associating every detected bounding box, the AssocA metric directly reflects the performance of AED. AED outperforms OVTrack by 5.8% (36.3%-38.4%) and 16.0% (33.2%-38.5%) on AssocA of base and novel categories of validation set respectively. For the test set, the increment for AssocA on base and novel categories are 11.9% (36.1%-40.4%) and 11.1% (26.2%-29.1%) respectively.
@@ IV-D1 SportsMOT
We further validate the CV-MOT performance of AED on the SportsMOT dataset. As shown in Tab II, for AED, we first use the weights trained on the $\mathcal{C}^{\mathrm{base}}$ of TAO following Section IV-C while the detector (YOLOX [38]) is trained on the training set of SportsMOT. Under such a setting, AED achieves a competitive HOTA score, i.e., 72.8, outperforming a large number of trackers [52, 24, 53, 5, 54, 18, 19, 55]. For a more equitable comparison, we also train AED, which is pretrained on COCO, on the training set of SportsMOT. AED achieves state-of-the-art performance on most metrics. Compared with Deep-EIoU [25], AED achieves the increment of 3.9% (74.1%-77.0%) on HOTA and 6.7% (75.0%-80.0%) on IDF1 by using only training set. Note that, unlike other tracking-by-detection methods leveraging some extra prior knowledge, AED only uses the similarities of features to achieve this performance.
@@ IV-E1 Analysis on Association-Centric Learning Mechanism
IV-E2 Analysis on Sampling Strategy
As the input of AED is a video clip in the training stage, there are different ways to merge a number of video frames into a clip. We conduct ablation experiments on two sampling methods on the TAO dataset as shown in Tab. VI. The first setting is that we fix the sampling interval to the number of three, i.e., skip two frames when sampling. Second, we allowed the sampling interval a to change randomly within the range of [1, 3], which introduces a wider variety of displacement and perspective changes during training. We find that the random sampling strategy improves the AssocA by 0.6 points.
TABLE VI: Different sampling strategy ablation on the validation set of TAO dataset.
Sampling Methods TETA LocA AssocA ClsA
fixed 54.9 71.7 51.5 41.6
random 55.2 71.8 52.1 41.7
IV-F Analysis on Parameter Sensitivity and Performance Ceiling
TABLE VII: Multi-head weight attention analysis on the validation set of TAO dataset.
@@ IV-F Analysis on Parameter Sensitivity and Performance Ceiling
IV-F1 Weight Attention Analysis
In Tab. VII, we conduct several experiments to validate the impact of different $h$ in the multi-head weight attention. We select several different values of $h$, i.e., 1, 2, 4, 8, and 16. It can be observed from Tab. VII that when $h$ equals 2, the model achieves a comparatively high performance.
IV-F2 Analysis on IoU Threshold to Filter Out Ambiguous Queries
As introduced in Section III-D, after the refinement in training, we calculate IoUs between boxes of object queries and their GTs to filter out the ambiguous ones, i.e., IoU $<\tau_{\mathrm{iou}}$. In order to investigate the impact of $\tau_{\mathrm{iou}}$ on AED, we conduct a comparative experiment in Tab. VIII. We find that a higher $\tau_{\mathrm{iou}}$ has a slight negative impact on the association performance (AssocA) of AED. Ultimately, we chose $\tau_{\mathrm{iou}}=0.5$ during training.
IV-F3 Analysis on the Length of Clips
During training, we take sampled clips of $n$ frames as the input. We conduct an experiment on the TAO dataset to find the most suitable $n$ in Tab. IX. We can find that with the increment of $n$, the performance of AED also improves. However, due to the limitation of GPU memory, $n=5$ represents the current upper limit of our experimental conditions.
IV-F4 Analysis on Performance Ceiling
TABLE VIII: $\tau_{iou}$ analysis on the validation set of TAO dataset.
@@ IV-F4 Analysis on Performance Ceiling
IV-G Visualization
IV-G1 Association-Centric Learning
TABLE IX: The impact of different clip lengths on the performance of AED.
@@ IV-G1 Association-Centric Learning
IV-G2 Tracking Results
In Fig. 6, we visualize some successful and failed tracking results of AED on different datasets including TAO, SportsMOT, and DanceTrack. Specifically, we use Co-DETR in TAO and YOLOX in SportsMOT and DanceTrack. AED is able to handle a wide range of scenarios including low frame rates (TAO), large movements(SportsMOT), severe occlusions (DanceTrack), and similar appearances (SportsMOT and DanceTrack). However, AED makes some mistakes in cases involving tiny objects (ID10, 12, etc of the first row of the TAO dataset) and highly dense scenes (ID 64, 65, etc of the second row of the TAO dataset).